3. Suppose that when the price of bottled water rises from $1.25 to $1.35 the daily quantity demanded falls from 5,000 to 4,000 bottles.

a. compute the elasticity of demand for bottled water.

b. interpret the number you calculated in part a.

c. if the price of bottle water were to rise by 5%, what would happen to the quantity demanded.
Answer #1

Ans) Elasticity is the responsiveness of quantity demanded to change in price. If elasticity is more than 1, good is elastic. If elasticity is less than 1,good is inelastic and if it is 1, good is unit elastic.

Elasticity of demand carries a negative sign, which shows negative relationship between price and quantity demanded.

Elasticity = %change in quantity demanded ÷ %change in price

Med-point formula for elasticity of Demand | Elasticity= Qa-Q P2-pi Hela y Trata/2 (P +P.2/21 4000-5000 2 (4000+5000)/2 11:35

2) Since elasticity is 2.89, good is elastic in nature.

3) Elasticity = %change in quantity ÷ %change in price

-2.89 = % change in quantity ÷ 5%

% change in quantity demanded = -2.89×5% = -14.45%

That is, quantity demanded will decrease by 14.45%